OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4>

OpenStudy (michele_laino):

\[\Large \cos \theta = \frac{{u \cdot v}}{{\left| u \right|\left| v \right|}}\]

OpenStudy (michele_laino):

where u.v is the scalar product between u and v, furthermore |u| is the length of the vector u, similarly for |v|

OpenStudy (michele_laino):

hint: \[\Large \begin{gathered} \left| u \right| = \sqrt {{6^2} + {{\left( { - 1} \right)}^2}} = ... \hfill \\ \left| v \right| = \sqrt {{7^2} + {{\left( { - 4} \right)}^2}} = ... \hfill \\ u \cdot v = 6 \times 7 + \left( { - 1} \right) \times \left( { - 4} \right) = ... \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

So 0.2º?

OpenStudy (michele_laino):

I got this: \[\Large \cos \theta = \frac{{u \cdot v}}{{\left| u \right|\left| v \right|}} = \frac{{46}}{{\sqrt {37} \sqrt {65} }} \cong 20.3{\text{degrees}}\]

OpenStudy (michele_laino):

oops.. \[\Large \cos \theta = \frac{{u \cdot v}}{{\left| u \right|\left| v \right|}} = \frac{{46}}{{\sqrt {37} \sqrt {65} }} \cong 0.93799\]

OpenStudy (michele_laino):

so: \[\Large \theta = 20.3\;{\text{degrees}}\]

OpenStudy (anonymous):

Thanks!

OpenStudy (michele_laino):

:)